Tom, a calendar year taxpayer, informs you that during the year, he incurs expenditures of $40,000 that qualify for the incremental research activities credit. In addition, Tom’s research-credit base amount for the year is $32,800.

Tom is in the 24% tax bracket. Determine which approach to the research expenditures and the research activities credit (other than capitalization and subsequent amortization) would provide the greater tax benefit.
Answer #1

Tom's incremental research activities credit for the year

Qualified expenditures for the research activities of the year = $40,000

Base Amount = $32,800

Incremental Research Expenditures = $40,000 - $32,800 = $7,200

Tax credit rate = 20%

Incremental Research Activities Credit = $7,200 * 20% = $1,440

Approach to the research expenditures and the research activities credit that would provide the greater tax benefit are:

Choice 1 - Reduce the deduction by 100% of credit and claim the full credit.

Qualified Expenditures - Credit = $40,000 - $1,440 = $38,560

Tax rate = 24%

Tax benefit of reduced deduction = $38,560 * 24% = $9,254.4

Allowed Credit = $1,440

Total Tax benefits of Choice 1 = $9,254.4 + $1,440 = $10,694.4

Choice 2 - Claim the full deduction and reduce the credit by 21%( the maximum corporate tax rate)

Qualified Expenditure = $40,000

Tax rate = 24%

Tax benefit of full deduction = $40,000 * 24% = $9,600

Reduced Credit = $1,440 - ($1,440 * 21%) = $1,137.6

Total Tax benefits of Choice 2 = $9,600 + $1,137.6 = $10,737.6

Thus, Choice 2 provides Tom a greater tax benefit of $43.2 ($10,737.6 - $10,694.4).